# Functional Central Limit Theorem for the simultaneous subgraph count of dynamic Erdős-Rényi random graphs

### v1

@@ Appendix A Auxiliary proofs
$=\mathcal{E}_{N}({\tt G}^{\star})\sqrt{\frac{\mathcal{F}_{N}\biggl{(}\bigl{(}\cup_{i\in{\tt V}({\tt G}^{\star})}g_{i}\bigr{)}\cap g_{j}\biggr{)}}{\mathcal{F}_{N}^{\rm{opt}}({H}_{f_{j}})}}\sqrt{\mathcal{F}_{N}\biggl{(}\bigl{(}\cup_{i\in{\tt V}({\tt G}^{\star})}g_{i}\bigr{)}\cap g_{j}\biggr{)}}$
Using the induction hypothesis (A.28) for $\mathcal{E}_{N}({\tt g}^{\star})$, from (3.10) (as $\bigl{(}\cup_{i\in{\tt V}({\tt G}^{\star})}g_{i}\bigr{)}\cap g_{j}{\,\sqsubset\,}g_{j}\sim{H}_{f_{j}}$), we obtain
@@
Proof of Lemma 7.2. Let us fix ${H}{\,\sqsubset\,}{\tt M}_{z}$ and $\boldsymbol{g}^{\cap}={(g^{\cap}_{1},\ldots,g^{\cap}_{z/2})}\in{\tt OCS}_{\boldsymbol{f}}({H})$. To calculate the cardinality $|\,{\tt G}_{\boldsymbol{f}}^{\star\star}({H},\boldsymbol{g}^{\cap})\,|$, we separate
@@
ways. Upon combining (A.30) and (A.31), the claim of Lemma 7.2 follows $\Box$

### v2

@@ Appendix A Auxiliary proofs
$=\mathcal{E}_{N}({\tt G}^{\star})\sqrt{\frac{\mathcal{F}_{N}\biggl(\bigl(\cup_{i\in{\tt V}({\tt G}^{\star})}g_{i}\bigr)\cap g_{j}\biggr)}{\mathcal{F}_{N}^{\rm{opt}}({H}_{f_{j}})}}\sqrt{\mathcal{F}_{N}\biggl(\bigl(\cup_{i\in{\tt V}({\tt G}^{\star})}g_{i}\bigr)\cap g_{j}\biggr)}{.}$
Using the induction hypothesis (A.28) for $\mathcal{E}_{N}({\tt g}^{\star})$, from (3.9) (as $\bigl(\cup_{i\in{\tt V}({\tt G}^{\star})}g_{i}\bigr)\cap g_{j}{\,{\subset}\,}g_{j}\sim{H}_{f_{j}}$), we obtain
@@
Proof of Lemma 8. Let us fix ${H}{\,{\subset}\,}{\tt M}_{z}$ and $\boldsymbol{g}^{\cap}={(g^{\cap}_{1},\ldots,g^{\cap}_{z/2})}\in{\tt OCS}_{\boldsymbol{f}}({H})$. To calculate the cardinality $|\,{\tt G}_{\boldsymbol{f}}^{\star\star}({H},\boldsymbol{g}^{\cap})\,|$, we separate
@@
ways. Upon combining (A.30) and (A.31), the claim of Lemma 8 follows. $\Box$